Use an appropriate local linear approximation to estimate the value of (122)^2/3
1
Expert's answer
2020-09-01T12:39:20-0400
"f(x)=x^{\frac {2}{3}}"

"f(x_0+\varDelta x)\approx f(x_0)+d[f(x_0)]"


"x_0=125, \varDelta x=-3"


"d[f(x_0)]=f^/(x_0)\times \varDelta x"


"f^/(x)=\frac {2}{3x^{\frac{1}{3}}}"


"d[f(x_0)]=f^/(x_0)\times \varDelta x=\frac {2}{15} \times(-3)=-0.4"


"f(x_0+\varDelta x)\approx f(x_0)+d[f(x_0)] \\ \approx 25-0.4=24.6"